# Dynamics of “Spontaneous” Topic Changes in Next Token Prediction with Self-Attention

### v1

@@ Abstract
Human cognition can spontaneously shift conversation topics, often triggered by emotional or contextual signals. In contrast, self-attention-based language models depend on structured statistical cues from input tokens for next-token prediction, lacking this spontaneity. Motivated by this distinction, we investigate the factors that influence the next-token prediction to change the topic of the input sequence. We define concepts of topic continuity, ambiguous sequences, and change of topic, based on defining a topic as a set of token priority graphs (TPGs). Using a simplified single-layer self-attention architecture, we derive analytical characterizations of topic changes. Specifically, we demonstrate that (1) the model maintains the priority order of tokens related to the input topic, (2) a topic change occurs only if lower-priority tokens outnumber all higher-priority tokens of the input topic, and (3) unlike human cognition, longer context lengths and overlapping topics reduce the likelihood of spontaneous redirection. These insights highlight differences between human cognition and self-attention-based models in navigating topic changes and underscore the challenges in designing conversational AI capable of handling “spontaneous” conversations more naturally. To our knowledge, this is the first work to address these questions in such close relation to human conversation and thought.
@@ 1 Introduction
In human cognition, topic changes often occur sporadically, with little apparent structure or forethought: an spontaneous shift in focus during a conversation, a sudden leap between ideas when brainstorming, or an unexpected redirection in storytelling. These seemingly chaotic shifts are intrinsic to the human way of processing and communicating information (Christoff and Fox,, 2018; Mills et al.,, 2020). For instance, consider a casual conversation where someone is discussing their favorite book, and then they suddenly mention an unrelated memory of a family gathering. This abrupt change may be due to an emotional connection, such as recalling reading that book during a family vacation, where sensory details like the scent of the ocean or the warmth of the sun trigger a vivid memory. The hippocampus helps link these sensory cues to specific memories, resulting in an unexpected yet meaningful shift (Lisman et al.,, 2017). The prefrontal cortex then integrates these memories, emotions, and contextual information, allowing for a spontaneous redirection of conversation (Miller and Cohen,, 2001).
Self-attention mechanisms in language models rely on contextual signals in the input to perform topic shifts, with next-token prediction guided by the relevance of preceding tokens (Vaswani et al.,, 2017). Unlike humans, who transition fluidly and intuitively between topics, often driven by emotion or creativity, transformers follow a structured, statistical approach. They remain on topic unless explicit cues signal a change, reflecting a key distinction between the somehow rigid processing of language models and the dynamic transitions of human cognition (Brown et al.,, 2020; Clark et al.,, 2019; Devlin et al.,, 2019; Bisk et al.,, 2020).
@@
Recent advancements in the related field have substantially deepened our understanding of self-attention mechanisms. The Transformer model, introduced by Vaswani et al., (2017), revolutionized natural language processing by using self-attention to determine the importance of each token in relation to others. This innovation enables Transformers to focus on relevant context when predicting the next token, significantly enhancing capabilities natural language understanding and generation (Brown et al.,, 2020; Devlin et al.,, 2019). Efforts to improve next token prediction have prioritized optimizing the efficiency of the self-attention mechanism, achieving empirical success across tasks such as machine translation, question answering, and text summarization. Additionally, connections between self-attention and support vector machines (SVM) have offered valuable insights into optimization strategies, and mechanisms within the next token prediction (Ataee Tarzanagh et al.,, 2023; Tarzanagh et al.,, 2023; Li et al., 2024b, ). These advancements underscore the transformative role of attention-based architectures in token prediction tasks. A recent study by Li et al., (2023) highlights that in mixed topic inputs, transformers achieve higher pairwise attention between same-topic words compared to different-topic words. However, our approach aligns more closely with human cognitive processes, where the input is typically centered on a single topic, ensuring that discussions or conversations remain within a consistent context or theme.
Despite these advancements, our understanding of how LLMs manage topic changes in comparison to human cognition remains limited. Human cognition seamlessly navigates such transitions through spontaneous thought and contextual awareness (Christoff and Fox,, 2018). Investigating this distinction could provide valuable insights into the strengths and limitations of language models in mimicking human-like conversation. Bridging this gap is essential for enhancing language models, particularly in their ability to adapt to and manage dynamic topic changes–a process closely aligned with the flow of spontaneous thought in human cognition. Specifically, we address the following research questions:
@@
To answer the above questions, it is essential first to explore how self-attention allocates attention to tokens for next token prediction with respect to the input topic. Understanding how attention weights are allocated offers key insights into the prioritization of tokens by self-attention mechanisms, particularly when models are trained on datasets spanning multiple topics. This fundamental understanding will pave the way for uncovering the role of attention patterns in driving “spontaneous” topic changes. To the best of our knowledge, this is the first work to explore these questions in direct relation with human conversation and thought.
@@
    Changing topics triggered by token frequency. In Theorem 4, we show that only if a lower-priority token appears more frequently than all higher-priority tokens of Topic A, the oracle’s responses may begin to reflect Topic B, signaling a shift.
@@
Figure 1: $\mathcal{G}^{(4)}_{a}$ and $\mathcal{G}^{(4)}_{b}$ (left) are token-priority graphs (TPGs) of the last token ${\bf e}_{4}$ for $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, respectively. $\mathcal{G}^{(4)}_{ab}$ (right) is the combination of TPG for $\mathbb{T}_{a}$ and TPG for $\mathbb{T}_{b}$.
Figure 2: Illustration of topic continuity, ambiguous sequence, and change of topic (see Definition 3,  4, and 5 for formal definitions). For each input sequence derived from $\mathbb{T}_{a}$, we employ a self-attention model trained on $\mathbb{T}_{a}$ and another model trained on the mixed-topics dataset $\mathbb{T}_{ab}$ to predict the next tokens $\hat{y}_{{\bf w}_{a}}$ and $\hat{y}_{{\bf w}_{ab}}$, respectively (orange). $\widehat{\mathcal{G}}^{(4)}_{ab}$ is the highest probability SCC (see Definition 1) in mixed-topics setting and $\widehat{\mathcal{G}}^{(4)}_{a}$ is the highest probability SCC in $\mathbb{T}_{a}$, The input sequence is classified as topic continuity if $\hat{y}_{{\bf w}_{ab}}\in\widehat{\mathcal{G}}^{(4)}_{a}(X)$ (top). The input sequence is ambiguous if $\hat{y}_{{\bf w}_{ab}}\not\in\widehat{\mathcal{G}}^{(4)}_{a}(X)$, but $\widehat{\mathcal{G}}^{(4)}_{a}(X)\subset\widehat{\mathcal{G}}^{(4)}_{ab}(X)$ (middle). ${\bf W}_{ab}$ is considered to change of topic, if $\hat{y}_{{\bf w}_{ab}}\not\in\widehat{\mathcal{G}}^{(4)}_{a}(X)$ and $\widehat{\mathcal{G}}^{(4)}_{a}(X)\not\subset\widehat{\mathcal{G}}^{(4)}_{ab}(X)$ (bottom). The numeric details are provided in Figure 6 in the appendix.
Here, we outline the structural overview of our approach. Following the framework of token-priority graphs (TPGs) introduced by Li et al., 2024b , we start with defining a topic as a set of TPGs (see Sec. 3). TPGs represent the priority order of tokens conditioned on the last token of the input sequence. For simplicity, we consider two different topics, Topic A and Topic B, characterized by two different sets of TPGs. We demonstrate the preservation of input topic priorities in a model trained with mixed-topic datasets (see Sec. 4). Additionally, we introduce three concepts to categorize the prediction outputs: topic continuity, ambiguous sequence, and change of topic. We subsequently establish the conditions under which a self-attention model induces a topic change, showing that the probability of topic changes does not increase with longer input sequences or the presence of overlapping topics (see Sec. 5). To formalize this approach, we begin with the structured problem setup.
@@ 2 Problem setup
Next topic prediction with self-attention model.
In line with the approach presented by Ataee Tarzanagh et al., (2023) and Li et al., 2024b , we frame the next-token prediction task as a multi-class classification problem. Given a vocabulary of size $K$ with an embedding matrix ${\bf E}=[{\bf e}_{1}\ {\bf e}_{2}\ \cdots\ {\bf e}_{K}]^{\top}\in\mathbb{R}^{K\times d}$, we aim to predict the next token ID $y\in[K]$ based on an input sequence ${\bf X}=[{\bf x}_{1}\ {\bf x}_{2}\ \cdots\ {\bf x}_{T}]^{\top}\in\mathbb{R}^{T\times d}$ with ${\bf x}_{i}\in{\bf E}$ for all $i\in[T]$. The training dataset, denoted as
@@
contains sequences of varying lengths $T_{i}$. In our notation ${\bf x}$ is the embedding vector corresponding to the token ID $x$, this is ${\bf x}={\bf e}_{x}$. For prediction, we utilize a single-layer self-attention model with a combined key-query weight matrix ${\bf W}\in\mathbb{R}^{d\times d}$ and identity value matrix as in Ataee Tarzanagh et al., (2023). The self-attention embedding output
@@
where $\mathbb{S}(\cdot)$ is the softmax operation and $\bar{{\bf x}}:={\bf x}_{T}$, serves as a weighted representation of the tokens, allowing for context-sensitive prediction of ${\bf y}$ based on the final input token. Let $\ell:\mathbb{R}\to\mathbb{R}$ be a loss function. For the training set DSET, we aim to minimize empirical risk (ERM) with:
@@
We keep the same first two assumptions from Li et al., 2024b :
@@
Assumption 1 represents a variation of the weight-tying approach commonly used in language models (Press and Wolf,, 2016; Vaswani et al.,, 2017). Once training is complete, for a new input sequence ${\bf X}$, and a model characterized by $\bf W$, we predict the next token ID $\hat{y}_{\bf w}$ based on greedy decoding the probabilities from the softmax of the classification output
@@
Li et al., 2024b defined a token-priority graph (TPG) as a directed graph with nodes representing tokens in the vocabulary. $\text{DSET}^{(k)}$ is a subset of sequences from DSET with the same last token is ${\bf e}_{k}=\bar{\bf x}$. They defined TPGs $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$ such that every $\mathcal{G}^{(k)}$ is a directed graph where for every sequence $({\bf X},{\bf y})\in\text{DSET}^{(k)}$ a directed edge is added from ${\bf y}$ to every token ${\bf x}\in{\bf X}$. TPGs are further divided into strongly-connected components (SCCs), which capture subsets of tokens where priority relations are cyclic, indicating comparable token importance. For tokens within two different SCCs, strict priority orders emerge, helping the model to differentiate between tokens when learning next-token predictions. We use the same notation as Li et al., 2024b , given a directed graph $\mathcal{G}$, for $i,j\in[K]$ such that $i\neq j$:
@@
Here a condensed version of the theorem:
Theorem 1 (from Li et al., 2024b ).
@@
This convergence implies that the model predicts the next token based on priorities obtained from the SCCs within the TPG relevant to the last token of the input sequence. We add here another reasonable assumption that prevents the probabilities in Equation 1 to be equal for improbable numerical reasons, and we present our first lemma.
@@
For any $({\bf X},y)\in\text{DSET}$, $\exists i,j\in[T]\text{ and }u,v\in\mathbb{Z}$ such that $u\left[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})\right]_{i}=v\left[\mathbb{S}({\bf X}{\tilde{\bf W}}\bar{{\bf x}})\right]_{j}\iff u=v\text{ and }\left[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})\right]_{i}=\left[\mathbb{S}({\bf X}{\tilde{\bf W}}\bar{{\bf x}})\right]_{j}$.
@@
Suppose conditions from Theorem 1 and Assumption 3 hold. Consider an input sequence ${\bf X}$ from DSET(k) and corresponding TPG $\mathcal{G}^{(k)}$, $\forall i,j\in[K]$ we have $\left[\mathbb{S}\left({\bf C}f_{\bf W}({\bf X})\right)\right]_{i}=\left[\mathbb{S}\left({\bf C}f_{\tilde{\bf W}}({\bf X})\right)\right]_{j}\iff(x_{i}\asymp x_{j})\in\mathcal{G}^{(k)}$.
@@ 3 Defining Topics
This assumption enables the application of the results from Li et al., 2024b , with the concepts of topics and TPGs being used interchangeably.
@@ 4 Attention within mixed topics
The attention priority similarity quantifies how well the weights ${\bf W}^{\prime}$ preserve the attention priority of the weights ${\bf W}$. A value of 1 indicates that the priority is fully preserved. Using this metric, we conduct experiments, with results in Figure 3. We generate embeddings with $K=10$ and $d=16$, and randomly construct TPGs for $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$. Using these TPGs, we randomly generate $\text{DSET}_{a}$ and $\text{DSET}_{b}$. We compute $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$, $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$ and $\frac{{\bf W}_{ab}(\tau)}{\|{\bf W}_{ab}(\tau)\|_{F}}$ using the same procedure as Li et al., 2024b . We generate test sequences ${\bf Z}$ within $\mathbb{T}_{a}$, and we calculate the attention priority similarity of $\frac{{\bf W}_{ab}(\tau)}{\|{\bf W}_{ab}(\tau)\|_{F}}$ relative to both $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ and $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$. We repeat this process for multiple TPGs and input sequences (details in Appendix). Figure 3 clearly demonstrates that the similarity converges to 1 after $\tau=8000$ iterations when evaluated relative to $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ (blue line), but fails to converge relative to $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$ (orange line). These observations align with the results of Theorem 3.

### v2

@@ Abstract
Human cognition can spontaneously shift conversation topics, often triggered by emotional or contextual signals. In contrast, self-attention-based language models depend on structured statistical cues from input tokens for next-token prediction, lacking this spontaneity. Motivated by this distinction, we investigate the factors that influence the next-token prediction to change the topic of the input sequence. We define concepts of topic continuity, ambiguous sequences, and change of topic, based on defining a topic as a set of token priority graphs (TPGs). Using a simplified single-layer self-attention architecture, we derive analytical characterizations of topic changes. Specifically, we demonstrate that (1) the model maintains the priority order of tokens related to the input topic, (2) a topic change can occur only if lower-priority tokens outnumber all higher-priority tokens of the input topic, and (3) unlike human cognition, longer context lengths and overlapping topics reduce the likelihood of spontaneous redirection. These insights highlight differences between human cognition and self-attention-based models in navigating topic changes and underscore the challenges in designing conversational AI capable of handling “spontaneous” conversations more naturally. To the best of our knowledge, no prior work has explored these questions with a focus as closely aligned to human conversation and thought.
@@ 1 Introduction
In human cognition, topic changes often occur sporadically, with little apparent structure or forethought: an spontaneous shift in focus during a conversation, a sudden leap between ideas when brainstorming, or an unexpected redirection in storytelling. These seemingly chaotic shifts are intrinsic to the human way of processing and communicating information (Christoff and Fox,, 2018; Mills et al.,, 2020). For instance, someone discussing their favorite book might suddenly mention a family gathering. This abrupt change may be due to an emotional connection, such as recalling reading that book during a family vacation, where sensory details like the scent of the ocean or the warmth of the sun trigger a vivid memory. The hippocampus helps link these sensory cues to memories, enabling a meaningful shift (Lisman et al.,, 2017). The prefrontal cortex then integrates these memories, emotions, and context to redirect conversations spontaneously (Miller and Cohen,, 2001).
However, large language models (LLMs) rely on contextual signals in the input to perform topic shifts. Unlike humans, who transition intuitively between topics, often driven by emotion or creativity, LLMs follow a structured, statistical approach, remaining on topic unless explicit cues signal a change. Figure 2 illustrates this distinction using the first sentence of the book “One hundred years of solitude” (García Márquez,, 1967). Additionally, related studies on conversational agents have recognized the practical challenges of spontaneous topic changes and proposed various practical approaches (Lim et al.,, 2010; Hwang et al.,, 2024; Xie et al.,, 2021; Soni et al.,, 2022; Ni et al.,, 2022; Lin et al.,, 2023).
@@
Recent advancements in the related field have substantially deepened our understanding of self-attention mechanisms in LLMs. The Transformer model, introduced by Vaswani et al., (2017), revolutionized natural language processing by using self-attention to determine the importance of each token in relation to others. This innovation enables Transformers to focus on relevant context when predicting the next token, significantly enhancing capabilities natural language understanding and generation(Brown et al.,, 2020; Clark et al.,, 2019; Devlin et al.,, 2019; Bisk et al.,, 2020). Additionally, connections between self-attention and support vector machines (SVMs) have offered valuable insights into optimization strategies, and mechanisms within the next token prediction (Tarzanagh et al., 2023b, ; Tarzanagh et al., 2023a, ; Li et al., 2024b, ). These advancements underscore the transformative role of attention-based architectures in token prediction tasks. A recent study by Li et al., (2023) highlights that in mixed topic inputs, transformers achieve higher pairwise attention between same-topic words compared to different-topic words. However, our approach aligns more closely with human cognitive processes, where the input is typically centered on a single topic, ensuring that discussions or conversations remain within a consistent context or theme.
Figure 1: Illustration of the difference between human cognition and AI contextual completion. The original fragment of “One hundred years of solitude" (García Márquez,, 1967) (top) has a clear “spontaneous" thought, but the GPT-2’s completion [^2] Just to illustrate, we use the prompt Please continue this short sentence, forgetting about “One hundred Years of Solitude”, since on a real conversation the LLM would be blind to the final output. (bottom), demonstrates continuity.
Despite these advancements, our understanding of how LLMs manage topic changes compared to human cognition remains limited. Human cognition seamlessly navigates such transitions through spontaneous thought and contextual awareness (Christoff and Fox,, 2018). Investigating this distinction could provide valuable insights into the strengths and limitations of language models in mimicking human-like conversation. Bridging this gap is essential for enhancing language models, particularly in their ability to adapt to and manage dynamic topic changes–a process closely aligned with the flow of spontaneous thought in human cognition. Specifically, we address the following research questions:
@@
To answer the above questions, it is essential first to explore how self-attention allocates attention to tokens for next token prediction with respect to the input topic. Understanding how attention weights are allocated offers key insights into the prioritization of tokens by self-attention mechanisms, particularly when models are trained on datasets spanning multiple topics. This fundamental understanding will pave the way for uncovering the role of attention patterns in driving “spontaneous” topic changes. To the best of our knowledge, there are no other studies that have investigated these dynamics so closely in relation to human conversation and thought processes.
@@
    Changing topics triggered by token frequency. In Theorem 4, we show that only if a lower-priority token appears more frequently than all higher-priority tokens of Topic A, the oracle’s responses may reflect a change of topic.
@@
Figure 2: Overview of our approach. We define a topic by a set of TPGs $\{\mathcal{G}^{(k)}\}^{K}_{k=1}$ (Def. 2) and generate a dataset for each topic such that the dataset can reconstruct the corresponding TPGs exactly (Assumption 4). The combination of dataset A and dataset B becomes the dataset for the mixed-topic model. We train self-attention models independently on each dataset. Then, we generate a new input sequence from topic A and predict the next token with two models, self-attention model A and self-attention model (mixed-topic). The next-token prediction with the mixed-topic model is categorized into three outcomes: keeps topic A (topic continuity from Def. 3), ambiguous sequence (from Def. 4), changes topic A (change of topic from Def. 5). Further details for each category are in Figure 4.
Following the framework of token-priority graphs (TPGs) introduced by Li et al., 2024b , we start with defining a topic as a set of TPGs (see Sec 3). For simplicity, we consider two different topics, Topic A and Topic B, characterized by two different sets of TPGs. We demonstrate the preservation of input topic priorities in a model trained with mixed-topic datasets (see Sec 4). Additionally, we introduce three concepts to categorize the prediction outputs: topic continuity, ambiguous sequence, and change of topic. Figure 2 outlines the structural overview of our approach. We subsequently establish the conditions under which a self-attention model induces a topic change, showing that the probability of topic changes does not increase with longer input sequences or the presence of overlapping topics (see Sec 5). All proofs are refer to Appendix A. To formalize this approach, we begin with the structured problem setup.
@@ 2 Problem Setup
Next Topic Prediction with Self-attention Model
In line with the approach presented by Tarzanagh et al., 2023b and Li et al., 2024b , we frame the next-token prediction task as a multi-class classification problem. Given a vocabulary of size $K$ with an embedding matrix ${\bf E}=[{\bf e}_{1}\ {\bf e}_{2}\ \cdots\ {\bf e}_{K}]^{\top}\in\mathbb{R}^{K\times d}$, we aim to predict the next token ID $y\in[K]$ based on an input sequence ${\bf X}=[{\bf x}_{1}\ {\bf x}_{2}\ \cdots\ {\bf x}_{T}]^{\top}\in\mathbb{R}^{T\times d}$ with ${\bf x}_{i}\in{\bf E}$ for all $i\in[T]$. The training dataset, denoted as
@@
contains sequences of varying lengths $T_{i}$. In our notation ${\bf x}$ is the embedding vector corresponding to the token ID $x$, this is ${\bf x}={\bf e}_{x}$. For prediction, we utilize a single-layer self-attention model with a combined key-query weight matrix ${\bf W}\in\mathbb{R}^{d\times d}$ and identity value matrix as in Tarzanagh et al., 2023b . The self-attention embedding output
@@
where $\mathbb{S}(\cdot)$ is the softmax operation and $\bar{{\bf x}}:={\bf x}_{T}$, serves as a weighted representation of the tokens, allowing for context-sensitive prediction of $y$ based on the final input token. Let $\ell:\mathbb{R}\to\mathbb{R}$ be a loss function. For the training dataset DSET, we consider the empirical risk minimization (ERM) with:
@@
We keep the first two assumptions from Li et al., 2024b :
@@
Assumption 1 represents a variation of the weight-tying approach commonly used in language models (Press and Wolf,, 2017; Vaswani et al.,, 2017). Once training is complete, for a new input sequence ${\bf X}$, and a model characterized by $\bf W$, we predict the next token ID $\hat{y}_{\bf w}$ based on greedy decoding the probabilities from the softmax of the classification output
@@
Li et al., 2024b defined a token-priority graph (TPG) as a directed graph with nodes representing tokens in the vocabulary. $\text{DSET}^{(k)}$ is a subset of sequences from DSET with the same last token is ${\bf e}_{k}=\bar{\bf x}$. They defined TPGs $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$ such that every $\mathcal{G}^{(k)}$ is a directed graph where for every sequence $({\bf X},y)\in\text{DSET}^{(k)}$ a directed edge is added from ${\bf e}_{y}$ to every token ${\bf x}\in{\bf X}$. TPGs are further divided into strongly-connected components (SCCs), which capture subsets of tokens where priority relations are cyclic, indicating comparable token importance. For tokens within two different SCCs, strict priority orders emerge, helping the model to differentiate between tokens when learning next-token predictions. We use the same notation as Li et al., 2024b , given a directed graph $\mathcal{G}$, for $i,j\in[K]$ such that $i\neq j$:
@@
Here is a condensed version of the theorem:
Theorem 1 (Li et al., 2024b ).
@@
This convergence implies that the model predicts the next token based on priorities obtained from the SCCs within the TPG relevant to the last token of the input sequence. Unlike the work in Li et al., 2024b , which considers multiple loss functions in subsequent results, we focus exclusively on a log-loss function in this work, leaving the exploration of other loss functions for future research. We add here another reasonable assumption that prevents the probabilities in Equation 1 to be equal for improbable numerical reasons, and we present our first lemma.
@@
For any $({\bf X},y)\in\text{DSET}$, $\exists i,j\in[T]\text{ and }u,v\in\mathbb{Z}$ such that $u\left[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})\right]_{i}=v\left[\mathbb{S}({\bf X}{\tilde{\bf W}}\bar{{\bf x}})\right]_{j}\text{ if and only if }u=v\text{ and }\left[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})\right]_{i}=\left[\mathbb{S}({\bf X}{\tilde{\bf W}}\bar{{\bf x}})\right]_{j}$.
@@
Suppose conditions from Theorem 1 and Assumption 3 hold. Consider an input sequence ${\bf X}$ from DSET(k) and corresponding TPG $\mathcal{G}^{(k)}$, $\forall i,j\in[K]$ we have $\left[\mathbb{S}\left({\bf C}f_{\bf W}({\bf X})\right)\right]_{i}=\left[\mathbb{S}\left({\bf C}f_{\tilde{\bf W}}({\bf X})\right)\right]_{j}\text{ iff }(x_{i}\asymp x_{j})\in\mathcal{G}^{(k)}$.
@@ 3 Defining Topics
Detailed explanation is provided in Appendix B. This assumption enables the application of the results from Li et al., 2024b , with the concepts of topics and TPGs being used interchangeably.
@@ 4 Attention within Mixed Topics
Figure 3: Illustration of topic combination. $\mathcal{G}^{(4)}_{a}$ and $\mathcal{G}^{(4)}_{b}$ are token-priority graphs (TPGs) of the last token ${\bf e}_{4}$ for $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, respectively. $\mathcal{G}^{(4)}_{ab}$ combines the TPG for $\mathbb{T}_{a}$ and the TPG for $\mathbb{T}_{b}$.
@@
Figure 4: Illustration of topic continuity, ambiguous sequence, and change of topic (see Definition 3,  4, and 5 for formal definitions). For each input sequence derived from $\mathbb{T}_{a}$, we employ a self-attention model trained on $\mathbb{T}_{a}$ and another model trained on the mixed-topics dataset $\mathbb{T}_{ab}$ to predict the next tokens $\hat{y}_{{\bf w}_{a}}$ and $\hat{y}_{{\bf w}_{ab}}$, respectively (orange). $\widehat{\mathcal{G}}^{(4)}_{ab}$ is the highest probability SCC (see Definition 1) in mixed-topics setting and $\widehat{\mathcal{G}}^{(4)}_{a}$ is the highest probability SCC in $\mathbb{T}_{a}$. Top: The input sequence is classified as topic continuity if $\hat{y}_{{\bf w}_{ab}}\in\widehat{\mathcal{G}}^{(4)}_{a}(X)$. Middle: The input sequence is ambiguous if $\hat{y}_{{\bf w}_{ab}}\not\in\widehat{\mathcal{G}}^{(4)}_{a}(X)$, but $\widehat{\mathcal{G}}^{(4)}_{a}(X)\subset\widehat{\mathcal{G}}^{(4)}_{ab}(X)$. Bottom: ${\bf W}_{ab}$ is considered to change of topic, if $\hat{y}_{{\bf w}_{ab}}\not\in\widehat{\mathcal{G}}^{(4)}_{a}(X)$ and $\widehat{\mathcal{G}}^{(4)}_{a}(X)\not\subset\widehat{\mathcal{G}}^{(4)}_{ab}(X)$. The numeric details are provided in Appendix C.3.
@@
The attention priority similarity quantifies how well the weights ${\bf W}^{\prime}$ preserve the attention priority of the weights ${\bf W}$. A value of 1 indicates that the priority is fully preserved. Using this metric, we conduct experiments, with results in Figure 5. We generate embeddings with $K=10$ and $d=16$, and randomly construct TPGs for $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$. Using these TPGs, we randomly generate $\text{DSET}_{a}$ and $\text{DSET}_{b}$. We compute $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$, $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$ and $\frac{{\bf W}_{ab}(\tau)}{\|{\bf W}_{ab}(\tau)\|_{F}}$ using the same procedure as Li et al., 2024b . We generate test sequences ${\bf Z}$ within $\mathbb{T}_{a}$, and we calculate the attention priority similarity of $\frac{{\bf W}_{ab}(\tau)}{\|{\bf W}_{ab}(\tau)\|_{F}}$ relative to both $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ and $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$. We repeat this process for multiple TPGs and input sequences (simulation details in Appendix C). Figure 5 clearly demonstrates that the similarity converges to 1 after $\tau=8000$ iterations when evaluated relative to $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ (blue line), but fails to converge relative to $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$ (orange line). These observations align with the results of Theorem 3.